# Self-supervised visual learning in the low-data regime: a comparative evaluation

### v1

@@ 3.4 Employed datasets
A summary of the datasets employed in the conducted experimental study can be found in Table 5, excluding the 19 VTAB ones used for the robustness experiments (Section 3.2); details about them are available in [88]. In certain cases, adjustments needed to be made for the data to be incorporated in the experimental evaluation, mainly targeting to fit to the defined ‘low-data regime’. In particular, Places-100 was constructed from Places-365 by randomly sampling 500 images per class, instead of 400 as in [78], for a total of 50,000 images (instead of 40,000) that were then segmented into a training/validation/test set following a 70-15-15 split ratio [^9] In the sequel, this modified variant is implied wherever Places-100 is mentioned. For the case of ADE20K, whose official on-line repository contains only a training set and a test set, the former one was randomly split into a training and a validation subset using a 80:20 ratio. Similarly, the labeled images of STL-10 were grouped into 10 predefined folds, each comprising 50 images from the total of 500 available annotated ones; then, one fold was utilized as the validation set and the remaining 9 formed the training one. Moreover, the SIXRay-100 dataset was resized to better fit the low-data regime, by constructing a subset of 300,000 images in the following way: all 8,929 positive images were retained and the negative ones were randomly sampled from the original negative images of the dataset; the formulated dataset was segmented following a 70-15-15 split ratio, resulting in a training/validation/test set of 210,000/45,000/45,000 images, respectively [^10] In the sequel, this modified low-data variant is implied wherever SIXRay-100 is mentioned. Overall, SIXRay-10 and the subsampled SIXRay-100 overlap in content, but the test images of SIXRay-10 are deliberately not included in the training or the validation set of SIXRay-100.
@@ 4 Experimental Results and Insights
Table 7: Top-5 accuracy in downstream classification for transfer-learning SSL setup (top-1 accuracy reported for STL-10 dataset). The ‘Sup-Large’ and DINOv2 baselines have been pretrained on large-scale datasets.
@@ 4.2 Robustness experimental results
This section discusses the outcomes obtained by the execution of the robustness evaluation protocols detailed in Section 3.2. In particular, Tables 9 and 10 illustrate downstream evaluation results for the Noisy-ImageNet-100, and the Imbalanced-ImageNet-100 and VTAB benchmarks, respectively. It must be highlighted that no downstream finetuning on Noisy-ImageNet-100 is performed, according to the relevant protocols. To this end, all SSL-pretrained models have been finetuned on regular ImageNet-100 and then evaluated on the three Noisy-ImageNet-100 datasets (in Table 9). On the other hand, the non-SSL baselines were trained from scratch on ImageNet-100 (‘Sup.’) or ImageNet-1k (‘Sup.-Large’) and then evaluated on Noisy-ImageNet-100. The DINOv2 baseline was pretrained on LVD-142M and then evaluated on Noisy-ImageNet-100, after appending to it a linear classifier finetuned on ImageNet-100. For the Imbalanced-ImageNet-100 benchmark (Table 10), the SSL methods and the basic supervised baselines were pretrained on ImageNet-100 and then finetuned on Imbalanced-ImageNet-100, before evaluation on the latter one’s test set. The ‘Sup.-Large’ baselines were pretrained on ImageNet-1k, while the off-the-shelf DINOv2 baseline was pretrained on LVD-142M. Moreover, the robustness evaluation protocol for VTAB (Table 10) is similar to the one for Imbalanced-ImageNet-100.
@@
From the results presented in Tables 9 and 10, it can be observed that MAE and DINO (both on ViT architectures) are proven to be the most robust low-data SSL approaches. However, no low-data SSL method is able to surpass in performance DINOv2 or the supervised pretraining/training from scratch baselines; not even the low-data baselines. The latter observation suggests that SSL pretraining in the low-data regime seems to be at a disadvantage with respect to robustness, but the reason is likely different among the three sets of experiments. In particular, the low-data supervised baselines are significantly less robust than their large-scale counterparts in the case of Noisy-ImageNet-100 and Imbalanced-ImageNet-100, but this is not true for VTAB. The particular characteristic of VTAB is that it uses a transfer learning setup with very few downstream finetuning examples, for evaluating how generally applicable different learnt representations are. This differs from both the transfer-learning setup, where there is a reasonably-sized downstream finetuning dataset available, and from the semi-supervised one, where the finetuning dataset simply has noisy ground-truth annotation and no transfer learning takes place. In this challenging scenario, MAE exhibits an advantage over DINO, under low-data pretraining conditions. Given that both methods are evaluated on ViT architectures, this finding suggests that MAE simply needs less pretraining data than DINO, in order to learn generally applicable features that are inherently semantically discriminant. It is likely that this behaviour is concealed in the transfer/semi-supervised-learning setup due to the larger finetuning dataset and the lack of transfer, respectively. Still, the low-data regime prevents even MAE from surpassing the supervised pretraining baselines.
@@ 4.3 Domain-specific experimental results
Method AID ChestX-Det SIXRay-10
SimCLR (ResNet) 58.1% 68.2% 73.3%
DINO (ViT) 65.7% 76.4% 74.1%
MAE (ViT) 63.5% 72.1% 69.7%
DeepClusterV2 (ResNet) 57.4% 70.6% 66.3%
@@
Sup.-Large (IN-1k, ResNet) 66.2% 72.5% 70.2%
Sup. (IN-100, ViT) 65.1% 73.6% 72.5%
Sup.-Large (IN-1k, ViT) 68.7% 74.5% 73.9%
DINOv2 (LVD-142M, ViT) 69.4 % 74.8% 73.5%
@@
This section presents in detail comparative evaluation results in domain-specific application cases, as detailed in Section 3.3. In particular, Table 11 illustrates downstream classification results for the AID (remote sensing), ChestX-Det (medical imaging) and SIXRay-10 (security imaging) datasets. AID is composed of aerial-view RGB images, while ChestX-Det and SIXRay-10 comprise X-ray scans. It needs to be highlighted that, although the remote sensing and the medical imaging experiments correspond to typical transfer-learning scenarios, the security imaging setting exhibits particular key characteristics: a) the pretraining and downstream datasets are overlapping in content, and b) there is extreme imbalance in the size of the supported classes (e.g., the ‘Negative’ class dominates the dataset, while the ‘Hammer’ one contains only 60 images). To this end, the security imaging experiment practically lies in-between the single-dataset and transfer-learning scenarios, while the downstream task is highly challenging, due to the increased class imbalance. Regarding experimental details, the MLRSNet, MedPix and SIXRay-100 datasets were used for SSL pretraining in the corresponding domains. In all three domain-specific experiments, SSL-pretrained models are compared against baselines pretrained in a supervised manner on ImageNet-100/ImageNet-1k, with the exception of DINOv2 that was pretrained using the LVD-142M dataset.
@@
From the results presented in Table 3.3, it can be observed that in the remote sensing domain (AID dataset) the large-scale pretraining baselines in general perform better than the low-data SSL approaches: pretrained DINOv2 and supervised ImageNet-1k pretraining lead to the highest performance. This suggests that, even though the remote sensing domain deviates from the one of natural images, the fact that it still remains in the RGB space leads the supervised pretraining methods to surpass the low-data SSL methods (of similar backbone architecture). However, in-domain low-data DINO pretraining does outperform low-data supervised pretraining on ImageNet-100. On the contrary, in the medical and security imaging domains in-domain low-data DINO pretraining outperforms all other approaches, including the large-scale pretraining ones. This is considered to stem from the significant irrelevance of (conventional) RGB supervised pretraining under a shift to the target X-ray domain. Moreover, the aforementioned difference in performance supports the intuition that in-domain low-data SSL pretraining is advantageous in highly domain-specific scenarios, compared to large-scale supervised pretraining schemes in conventional RGB settings. It is notable to mention that, given the extreme class imbalance of SIXRay, SSL methods are able to learn generic features from all images, including those of the dominant ‘Negative’ class, even under a low-data pretraining setting. In contrast, large-scale supervised pretraining on ImageNet-1k does not lead to modeling better representations (most likely due to feature space discrepancy) and the potential of the subsequent supervised finetuning is limited by the class imbalance issue.
@@ 4.4 Explanation of learnt representations
Table 14: Median silhouette coefficients for domain-specific downstream scenarios
@@
Method AID ChestX-Det SIXRay-10
SimCLR (ResNet) 0.44 0.47 0.53
DINO (ViT) 0.58 0.61 0.54
MAE (ViT) 0.51 0.48 0.60
DeepClusterV2 (ResNet) 0.45 0.50 0.48
@@
Sup. (ViT) 0.53 0.58 0.59
Sup.-Large (ResNet) 0.52 0.57 0.56
Sup.-Large (ViT) 0.57 0.60 0.61
DINOv2 (ViT) 0.59 0.56 0.63
@@
Tables 12, 13 and 14 illustrate the estimated median silhouette coefficient $S_{m}$ for the considered object-centric, scene-centric and domain-specific downstream scenarios, respectively, from Sections 4.1 and 4.3. It must be noted that for the single-dataset setup in Tables 12 and 13, ‘Sup.’ and ‘Sup.-Large’ both imply a single setting, i.e., no pretraining/random initialization. In all cases, the median $S_{m}$ for each model was calculated by obtaining its inferred representations for the test images of each downstream dataset, after finetuning. As can be seen, the obtained quantitative explanations are in general closely correlated with the respective downstream classification accuracy scores (reported in Sections 4.1 and 4.3), indicating that indeed better performing SSL methods owe their discrimination capability to their ability to form more distinct, compact and well-separated class representations. This observation also holds for the supervised pretraining baselines and for the DINOv2 baseline, which is pretrained on the giant LVD-142M dataset and whose high test accuracy in the majority of transfer-learning setup cases is linked to its ability to generate very accurate/robust class representations with low intra-class variance and high inter-class separation in a zero-shot way; this is also in line with the foundation model status of off-the-shelf DINOv2 models. Additionally, it needs to be highlighted that in-domain low-data SSL pretraining in general leads to almost equally good or even better downstream image representations in domain-specific experimental setting, as also observed in the respective downstream classification tasks.

### v2

@@ 3.4 Employed datasets
CEM500K [70]
It comprises 496,544 cellular electron microscopy images. The main dataset is not annotated, since it was constructed for enabling SSL pretraining.
RaabinWBC [74]
It contains 16,633 microscopy images of 5 classes of white blood cells, annotated for whole-image classification. The training/validation/test subset contains 10,185/4,339/2,119 images, respectively.
BIRDSAI [75]
It contains 61,994 real and 101,542 synthetic infrared images captured from a Unmanned Aerial Vehicle (UAV) view, during night-time surveillance of protected areas in Africa. The real images have been divided into two distinct sets: a training set comprising 40,661 images and a test set comprising 21,333 images. All synthetic images have been incorporated into the training set. It is annotated for object detection using 2 classes (‘Animals’, ‘Humans’).
HIT-UAV [76]
It contains 2,898 thermal infrared images extracted from UAV video footage in urban settings. It is divided into a training/validation/test subset, of 2,029/290/579 images, respectively, and is annotated for object detection using 4 classes (‘Person’, ‘Car’, ‘Bicycle’, ‘Other Vehicle’).
Table 4: Image datasets employed in the evaluation study (continued)
A summary of the datasets employed in the conducted experimental study can be found in Table LABEL:tab::Datasets, excluding the 19 VTAB ones used for the robustness experiments (Section 3.2); details about them are available in [66]. In certain cases, adjustments needed to be made for the data to be incorporated in the experimental evaluation, mainly targeting to fit to the defined ‘low-data regime’. In particular, Places-100 was constructed from Places-365 by randomly sampling 500 images per class, instead of 400 as in [56], for a total of 50,000 images (instead of 40,000) that were then segmented into a training/validation/test set following a 70-15-15 split ratio [^5] In the sequel, this modified variant is implied wherever Places-100 is mentioned. For the case of ADE20K, whose official on-line repository contains only a training set and a test set, the former one was randomly split into a training and a validation subset using a 80:20 ratio. Similarly, the labeled images of STL-10 were grouped into 10 predefined folds, each comprising 50 images from the total of 500 available annotated ones; then, one fold was utilized as the validation set and the remaining 9 formed the training one. Moreover, the SIXRay-100 dataset was resized to better fit the low-data regime, by constructing a subset of 300,000 images in the following way: all 8,929 positive images were retained and the negative ones were randomly sampled from the original negative images of the dataset; the formulated dataset was segmented following a 70-15-15 split ratio, resulting in a training/validation/test set of 210,000/45,000/45,000 images, respectively [^6] In the sequel, this modified low-data variant is implied wherever SIXRay-100 is mentioned. Overall, SIXRay-10 and the subsampled SIXRay-100 overlap in content, but the test images of SIXRay-10 are deliberately not included in the training or the validation set of SIXRay-100. Finally, the variant of the CEM500K dataset utilized in this study has been derived after random subsampling to 300,000 images, so that it falls under the defined low-data range, and segmentation into a training and a validation subset using a 80:20 ratio.
@@ 4 Experimental Results and Insights
Table 6: Top-5 accuracy in downstream classification for transfer-learning SSL setup (top-1 accuracy reported for STL-10 dataset). The ‘Sup.-Large’ and DINOv2 baselines have been pretrained on large-scale datasets.
@@ 4.1 Main experimental results
50505050656565658080808010010010010060606060656565657070707075757575IN-100 Sample Size (%)Downstream Accuracy (%)SimCLR (ResNet)None/Random (ResNet)
Figure 8: Top-5 accuracy in single-dataset downstream classification as a function of IN-100 sample size
@@ 4.2 Robustness experimental results
This section discusses the outcomes obtained by the execution of the robustness evaluation protocols detailed in Section 3.2. In particular, Tables 8 and 9 illustrate downstream evaluation results for the Noisy-ImageNet-100, and the Imbalanced-ImageNet-100 and VTAB benchmarks, respectively. It must be highlighted that no downstream finetuning on Noisy-ImageNet-100 is performed, according to the relevant protocols. To this end, all SSL-pretrained models have been finetuned on regular ImageNet-100 and then evaluated on the three Noisy-ImageNet-100 datasets (in Table 8). On the other hand, the non-SSL baselines were trained from scratch on ImageNet-100 (‘Sup.’) or ImageNet-1k (‘Sup.-Large’) and then evaluated on Noisy-ImageNet-100. The DINOv2 baseline was pretrained on LVD-142M and then evaluated on Noisy-ImageNet-100, after appending to it a linear classifier finetuned on ImageNet-100. For the Imbalanced-ImageNet-100 benchmark (Table 9), the SSL methods and the basic supervised baselines were pretrained on ImageNet-100 and then finetuned on Imbalanced-ImageNet-100, before evaluation on the latter one’s test set. The ‘Sup.-Large’ baselines were pretrained on ImageNet-1k, while the off-the-shelf DINOv2 baseline was pretrained on LVD-142M. Moreover, the robustness evaluation protocol for VTAB (Table 9) is similar to the one for Imbalanced-ImageNet-100.
@@
From the results presented in Tables 8 and 9, it can be observed that MAE and DINO (both on ViT architectures) are proven to be the most robust low-data SSL approaches. However, no low-data SSL method is able to surpass in performance DINOv2 or the supervised pretraining/training from scratch baselines; not even the low-data baselines. The latter observation suggests that SSL pretraining in the low-data regime seems to be at a disadvantage with respect to robustness, but the reason is likely different among the three sets of experiments. In particular, the low-data supervised baselines are significantly less robust than their large-scale counterparts in the case of Noisy-ImageNet-100 and Imbalanced-ImageNet-100, but this is not true for VTAB. The particular characteristic of VTAB is that it uses a transfer learning setup with very few downstream finetuning examples, for evaluating how generally applicable different learnt representations are. This differs from both the transfer-learning setup, where there is a reasonably-sized downstream finetuning dataset available, and from the semi-supervised one, where the finetuning dataset simply has noisy ground-truth annotation and no transfer learning takes place. In this challenging scenario, MAE exhibits an advantage over DINO, under low-data pretraining conditions. Given that both methods are evaluated on ViT architectures, this finding suggests that MAE simply needs less pretraining data than DINO, in order to learn generally applicable features that are inherently semantically discriminant. It is likely that this behaviour, which is compatible with the findings of [49], gets concealed in the transfer/semi-supervised-learning setup due to the larger finetuning dataset and the lack of transfer, respectively. Still, the low-data regime prevents even MAE from surpassing the supervised pretraining baselines.
@@ 4.3 Domain-specific experimental results
Method AID ChestX-Det SIXRay-10 RaabinWBC
SimCLR (ResNet) 58.1% 68.2% 73.3% 62.4%
DINO (ViT) 65.7% 76.4% 74.1% 80.5%
MAE (ViT) 63.5% 72.1% 69.7% 78.2%
DeepClusterV2 (ResNet) 57.4% 70.6% 66.3% 59.3%
@@
Sup.-Large (IN-1k, ResNet) 66.2% 72.5% 70.2% 42.4%
Sup. (IN-100, ViT) 65.1% 73.6% 72.5% 39.4%
Sup.-Large (IN-1k, ViT) 68.7% 74.5% 73.9% 34.5%
DINOv2 (LVD-142M, ViT) 69.4 % 74.8% 73.5% 85.0%
Table 11: Mean Average Precision in downstream object detection for the UAV-captured infrared domain-specific experiments, using the HIT-UAV downstream dataset
Metrics
Method mAP50 mAP50:95
SimCLR (ResNet) 0.804 0.497
DINO (ViT) 0.854 0.529
MAE (ViT) 0.837 0.518
DeepClusterV2 (ResNet) 0.812 0.500
[…]
Sup.-Large (IN-1k, ResNet) 0.760 0.456
Sup. (IN-100, ViT) 0.802 0.497
Sup.-Large (IN-1k, ViT) 0.810 0.502
DINOv2 (LVD-142M, ViT) 0.818 0.507
@@
This section presents in detail comparative evaluation results in domain-specific application cases, as detailed in Section 3.3. In particular, Table 10 illustrates downstream classification results for the AID (remote sensing), ChestX-Det (medical imaging), SIXRay-10 (security imaging) and RaabinWBC (microscopy imaging) datasets. AID is composed of aerial-view RGB images, ChestX-Det and SIXRay-10 comprise X-ray scans, while RaabinWBC contains microscope images of white blood cells. It needs to be highlighted that the security imaging setting exhibits particular key characteristics: a) the pretraining and downstream datasets are overlapping in content, and b) there is extreme imbalance in the size of the supported classes (e.g., the ‘Negative’ class dominates the dataset, while the ‘Hammer’ one contains only 60 images). To this end, the security imaging experiment practically lies in-between the single-dataset and transfer-learning scenarios, while the downstream task is highly challenging, due to the increased class imbalance. Regarding experimental details, the MLRSNet, MedPix, SIXRay-100 and CEM500K datasets were used for SSL pretraining in the corresponding domains. Finally, Table 11 depicts downstream object detection results for the HIT-UAV dataset (UAV-captured infrared domain), after pretraining on the BIRDSAI dataset. In all domain-specific experiments, both for whole-image classification and for object detection, SSL-pretrained models are compared against baselines pretrained in a supervised manner on ImageNet-100/ImageNet-1k, with the exception of DINOv2 that was pretrained using the LVD-142M dataset.
@@
From the experimental results, it can be observed that in the remote sensing domain (AID dataset) the large-scale pretraining baselines in general perform better than the low-data SSL approaches: pretrained DINOv2 and supervised ImageNet-1k pretraining lead to the highest performance. This suggests that, even though the remote sensing domain deviates from the one of natural images, the fact that it still remains in the RGB space leads the supervised pretraining methods to surpass the low-data SSL methods (of similar backbone architecture). However, in-domain low-data DINO pretraining does outperform low-data supervised pretraining on ImageNet-100. On the contrary, in the medical, security and UAV-captured infrared imaging domains in-domain low-data DINO pretraining outperforms all other approaches, including the large-scale pretraining ones. This is considered to stem from the significant irrelevance of (conventional) RGB supervised pretraining under a shift to the target domain. Moreover, the aforementioned difference in performance supports the intuition that in-domain low-data SSL pretraining is advantageous in highly domain-specific scenarios, compared to large-scale supervised pretraining schemes in conventional RGB settings. It is notable to mention that, given the extreme class imbalance of SIXRay, SSL methods are able to learn generic features from all images, including those of the dominant ‘Negative’ class, even under a low-data pretraining setting. In contrast, large-scale supervised pretraining on ImageNet-1k does not lead to modeling better representations (most likely due to feature space discrepancy) and the potential of the subsequent supervised finetuning is limited by the class imbalance issue. Finally, in the microscopy imaging domain, in-domain low-data DINO pretraining outperforms all other approaches, including the large-scale pretraining ones, except zero-shot DINOv2. A potential reason for this is that the gigantic LVD-142M pretraining dataset may indeed contain microscopy images, but since it has not been publicly released this remains a conjecture.
@@ 4.4 Explanation of learnt representations
Table 14: Median silhouette coefficients for domain-specific classification downstream scenarios
@@
Method AID ChestX-Det SIXRay-10 RaabinWBC
SimCLR (ResNet) 0.44 0.47 0.53 0.31
DINO (ViT) 0.58 0.61 0.54 0.37
MAE (ViT) 0.51 0.48 0.60 0.35
DeepClusterV2 (ResNet) 0.45 0.50 0.48 0.28
@@
Sup. (ViT) 0.53 0.58 0.59 0.24
Sup.-Large (ResNet) 0.52 0.57 0.56 0.26
Sup.-Large (ViT) 0.57 0.60 0.61 0.24
DINOv2 (ViT) 0.59 0.56 0.63 0.41
@@
Tables 12, 13 and 14 illustrate the estimated median silhouette coefficient $S_{m}$ for the considered object-centric, scene-centric and domain-specific classification downstream scenarios, respectively, from Sections 4.1 and 4.3. It must be noted that for the single-dataset setup in Tables 12 and 13, ‘Sup.’ and ‘Sup.-Large’ both imply a single setting, i.e., no pretraining/random initialization. In all cases, the median $S_{m}$ for each model was calculated by obtaining its inferred representations for the test images of each downstream dataset, after finetuning. As can be seen, the obtained quantitative explanations are in general closely correlated with the respective downstream classification accuracy scores (reported in Sections 4.1 and 4.3), indicating that indeed better performing SSL methods owe their discrimination capability to their ability to form more distinct, compact and well-separated class representations. This observation also holds for the supervised pretraining baselines and for the DINOv2 baseline, which is pretrained on the giant LVD-142M dataset and whose high test accuracy in the majority of transfer-learning setup cases is linked to its ability to generate very accurate/robust class representations with low intra-class variance and high inter-class separation in a zero-shot way; this is also in line with the foundation model status of off-the-shelf DINOv2 models. Additionally, it needs to be highlighted that in-domain low-data SSL pretraining in general leads to almost equally good or even better downstream image representations in domain-specific experimental setting, as also observed in the respective downstream classification tasks.
@@
4.5 Summary of Key Experimental Insights
20k40k60k80k100k120k140k160k180k200k220k240k55-5- 5005555101010101515151520202020252525253030303035353535404040404545454550505050ChestX-DetSUN DBAIDCOCOSIXRay-10RaabinWBCPretraining Set SizeDINO/ViT Accuracy Gain Over Sup./ViT (%)MedPix (32k)P-100 (35k)MLRSNet (80k)IN-100 (100k)SIXRay-100 (210k)CEM500K (240k)
Figure 10: DINO/ViT gains compared to Sup./ViT, in the transfer-learning experiments (best viewed in color)
20k40k60k80k100k120k140k160k180k200k220k240k1010-10- 1055-5- 5005555101010101515151520202020252525253030303035353535404040404545454550505050ChestX-DetSUN DBAIDCOCOSIXRay-10RaabinWBCPretraining Set SizeDINO/ViT Accuracy Gain Over Sup.-Large/ViT (%)MedPix (32k)P-100 (35k)MLRSNet (80k)IN-100 (100k)SIXRay-100 (210k)CEM500K (240k)
Figure 11: DINO/ViT gains compared to Sup.-Large/ViT in the transfer-learning experiments (best viewed in color)
This study showed that in the single-dataset setup, where cross-dataset generalization is not prioritized, CNN-based SSL methods (e.g., SimCLR on ResNet) outperform Vision Transformers and ViT-oriented SSL methods (e.g., DINO) in the low-data regime, potentially due to the data hunger of Vision Transformers. A notable insight is the identification of a pretraining dataset size threshold (estimated between 50k-65k images, assuming identical pretraining and downstream image domains), below which SSL becomes ineffective for learning useful representations. In the main transfer-learning setup, DINO on ViT architectures emerges as the leading SSL method within the low-data regime, though it does not surpass the performance of conventional supervised pretraining approaches.
In the semi-supervised single-dataset setup, DINO performs strongly, attributed to its capacity to learn more universally applicable embeddings that maintain effectiveness even in low-data pretraining. This ability proves especially advantageous in highly domain-specific transfer-learning settings, where in-domain low-data DINO pretraining consistently outperforms supervised large-scale pretraining on conventional RGB datasets in downstream accuracy. This advantage is most pronounced in fields significantly different from conventional natural image datasets, such as medical or security imaging
Representation analysis reveals that DINO’s superior discrimination ability compared to other SSL methods is due to its formation of distinct, compact, and well-separated class representations, with a broader focus across the input image, leading to lower uncertainty in feature generation. In contrast, MAE shows a focused attention on the visible foreground object, supporting higher robustness under low-data pretraining conditions and needing less pretraining data to learn semantically discriminative features. However, its robustness is still lower than that of both large-scale and low-data supervised pretraining approaches, particularly in noisy or imbalanced settings.
In order to more clearly showcase that any identified gains of low-data SSL pretraining over supervised pretraining in the transfer-learning setup are due to in-domain pretraining and not due to any differences in the size of the pretraining datasets, across the various sets of experiments, Figures 10 and 11 visualize part of the information contained in Tables 6 and 10. The horizontal axis is marked with the sizes of the SSL pretraining datasets, while the vertical axis represents downstream classification accuracy gains. Figure 10/11 depicts the reported DINO/ViT gains compared to Sup. (ViT)/Sup.-Large (ViT), respectively. As it can be seen, COCO and SUN DB accuracy scores were the ones selected to be visualized for SSL pretraining in IN-100 and P-100, due to the significantly larger test set sizes of these downstream datasets over STL-10 and ADE20K. In general, it is evident that any downstream classification accuracy gains of DINO do not fully correlate with pretraining dataset size, indicating that pretraining domain relevance is also important.